OpenStudy (anonymous):

A lottery has one $2000 prize, two $1000 prizes, and ten $500 prizes. Two thousand tickets are sold at $5 each. Find the expected value if a person buys two tickets.

OpenStudy (anonymous):

we do it the snap way

OpenStudy (anonymous):

buy up all the tickets spend \($5\times 2000=$10,000\) get back \(\$2000+2\times \$1000+10\times\$500=$9000\) for a net loss of \($10,000-\$9000=\$1000\)

OpenStudy (anonymous):

averages over the number of tickes we have lost \(\frac{\$1000}{2000}=.50\) i.e. we lost 50 cents a ticket. two tickes expected loss is $1 ie expacted values is -1

OpenStudy (anonymous):

thank you ;))